# No-exclaves percolation on random networks

### v1

@@ 3.4 The average NExP size and percolation threshold
We derive expressions for the average size of small NExP components, excluding the giant NExP component. When the giant NExP component does not exist, a node reached by following a randomly chosen link must belong to either a small component with probability $u$ or a giant unoccupied component with probability $1-w$, resulting in the condition $11-w+u=$ 1. It simplifies to the condition $u=w$. With this condition, the average small NExP component size can be expressed as
@@
$=F_{0}(1)+[qG_{0}^{\prime}(1)+(1-q)G_{0}^{\prime}(w)]F_{1}^{\prime}(1),$ (12)
where ${}^{\prime}$ denotes the derivative with respect to its argument $x$. From Eq. (7), we obtain
$F_{1}^{\prime}(1)=\frac{w}{1-qG_{1}^{\prime}(1)-(1-q)G_{1}^{\prime}(w)}.$ (13)
@@
$\chi=F_{0}(1)+\frac{w[qG_{0}^{\prime}(1)+(1-q)G_{0}^{\prime}(w)]}{[1-qG_{1}^{\prime}(1)-(1-q)G_{1}^{\prime}(w)]}.$ (14)
@@
$q_{c}G_{1}^{\prime}(1)+(1-q_{c})G_{1}^{\prime}(w)=1.$ (15)
The percolation threshold $q_{c}$ of the giant NExP component is in general not amenable to compact expression in terms of $z$ or other moments of $P(k)$ because $w$ also contains $q$, yet can be computed numerically. Beyond this threshold $q_{c}$, the size of the giant NExP component continues to grow as more occupied nodes are added.
In NExP, there is another transition where a giant unoccupied component disappears as $q$ increases. In a scenario where a giant unoccupied component vanishes, there is a potential for all nodes within a connected component to be part of the giant NExP component. We refer the point of this transition as the “complete” percolation point, denoted as $q^{*}$. The complete percolation point $q^{*}$ corresponds to the location in which the size $\Phi$ of the giant NExP component becomes unity if all nodes of the underlying network belong to a single connected component. The location of $q^{*}$ can be identified by the divergence of the average size $\langle s\rangle$ of unoccupied components. The condition for $q^{*}$ is given from Eq. (3) as
@@
which is directly related to the well-known result for the percolation on random networks callaway ; cohen2000 ; molloy1998 .
@@
The transitions can be interpreted by the stability analysis of the solutions of Eqs. (5) and (9). The complete percolation point $q^{*}$ corresponds to the point where the trivial fixed point of $w$ in Eq. (5) becomes unstable. The trivial solution of Eq. (5) is $w=1$ since $G_{1}(1)$ is normalized to be $1$. The solution $w=1$ corresponds to no giant unoccupied component considering the meaning of $w$. That means the giant unoccupied component disappeared at $q^{*}$. On the other hand, the NExP threshold $q_{c}$ can be obtained by the stability analysis of Eq. (9). The trivial solution of Eq. (9) is $u=w$ because Eq. (9) reduces into Eq. (5) when $u=w$, or equivalently $11-w+u=$ 1. The solution $u=w$ represents no giant NExP component and it becomes unstable at $q_{c}$. Therefore another solution becomes stable, leading to the emergence of the giant NExP component.
@@
These two transition points $q_{c}$ and $q^{*}$ do not in general coincide on random networks. The percolation transition $q_{c}$ is generally smaller than the complete point $q^{*}$. Therefore, there can be three phases in NExP, depending on the existence and extinction of the giant NExP component and the giant unoccupied component: We have phases in which i) the giant unoccupied component exists but giant NExP component does not when $q<q_{c}$; ii) both giant NExP component and giant unoccupied component exist when $q_{c}<q<q^{*}$; iii) the giant NExP component exists but giant unoccupied component does not when $q>q^{*}$. The multiple percolation transitions in network NExP are in stark contrast to the 2D NExP in which the two occur at the same point nexp . Our solutions suggest that the discontinuous percolation transition into complete percolation displayed by 2D NExP can be resolved into two transitions combined into one.
4 Results
@@ 4.1 On Random Regular Networks
When $w=1$ for $q>q^{*}$, $u$ becomes 00. If $q<q^{*}$ indicating that $w=q/(1-q)$, the possible solutions of $u$ are either $u=q/(1-q)$ and $(1-2q)^{2}/(1-q)$. The solution $u=w=q/(1-q)$ corresponds to the case when there is no giant NExP component. If $u=(1-2q)^{2}/(1-q)$, $1-w+u$ becomes less than unity meaning that there is a giant NExP component. The percolation threshold between two phases is given by
$w(q_{c})$ $=\frac{1}{z}=\frac{1}{3}.$ (24)
@@
Therefore we have three regimes of no-exclaves percolation in RR networks: i) a giant NExP component does not exist but a giant unoccupied component exists ($q<q_{c}$); ii) both giant NExP component and giant unoccupied component exist ($q_{c}<q<q^{*}$); and iii) an only giant NExP component whose size is unity exists ($q>q^{*}$). Since we can express $u$ and $w$ in closed forms for RR networks with $z=3$, we can explicitly obtain the expression for the size of the giant NExP component $\Phi$ as:
@@
As shown in Fig. 2, the probabilities $w$ and $u$, and size of the giant NExP component $\Phi$ as a function of the occupation probability $q$ show perfect agreement with the Monte Carlo simulation results.
Figure 2: Analytic predictions of $\Phi$, $w$, and $1-w+u$ with respect to $q$ on RR networks with $z=3$ and $N=10^{6}$ are shown. Theoretical results (lines) and numerical results (points) are shown together. Error-bars denote the standard deviation from $10^{6}$ independent runs.
@@ 4.2 On Erdős-Rényi networks
We consider the no-exclaves percolation model on Erdős-Rényi (ER) networks, which have a Poisson degree distribution $P(k)=e^{-z}z^{k}/k!$ in the large $N$ limit, where $z$ is the average degree. The generating functions for ER networks are given by $G_{0}(x)=e^{z(x-1)}$ and $G_{1}(x)=e^{z(x-1)}$. This leads us to the coupled equations for the variables $w$ and $u$, given by
@@
The complete percolation point and the percolation threshold for NExP can be determined by Eqs. (17) and (15), respectively. These conditions show the instability of the trivial solutions, $w=1$ for the complete percolation point and $u=w$ for the percolation threshold. For ER networks, the specific values for these thresholds are given by
$q^{*}=1-\frac{1}{z},\quad w(q_{c})=\frac{1}{z}.$ (36)
We have three regimes in the NExP model, similar to RR networks. i) When $q<q_{c}$, there is no giant NExP component but a giant unoccupied component exists. ii) When $q_{c}<q<q^{*}$, both a giant NExP component and a giant unoccupied component exist. iii) When $q>q^{*}$, only a giant NExP component exists. Furthermore, when $q>q_{c}$, the size of the giant NExP component $\Phi$ can be expressed as
$\Phi$ $=1-qe^{z(u-w)}-(1-q)[1-e^{z(w-1)}+e^{z(u-1)}].$ (37)
where $w$ and $u$ can be obtained by Eqs. (34-35).
We compare the theory derived using the generating function method with Monte Carlo simulations on ER networks with average degree $z=3/2,2,3$ and network size $N=10^{6}$. Figure 3 shows the size of the largest NExP component, $\Phi$, obtained from our theory (line) and numerical simulations (points) as a function of $q$. We observe excellent agreement between our theoretical predictions and the simulation results. Furthermore, we validate the locations of the percolation transition, $p_{c}$, and the complete transition point, $p^{*}$, which were predicted from Eq. (36).
Figure 3: The size of the giant NExP component $\Phi$ as a function of the occupation probability $q$ on ER networks with mean degree $z=3/2,2$, and $3$ and $N=10^{6}$. Numerical results (points) averaged over $10^{6}$ independent runs and analytic curves (lines) are shown together.
@@ 4.3 On real-world networks
We test our theory for the no-exclaves percolation on top of real-world networks built from empirical dataset. We use the two datasets: the Internet of autonomous system level caida and the U.S. airport network in 2010 obtained from “Open flights” flights . The router network of Internet represents that the traffic flow exchanges with its neighboring peers at autonomous system level, collected on November 5, 2007. The U.S. airport network represents the connection from the origin to destination airport. For the sake of simplicity, we ignore directionality and weights of the links from the data when we construct the networks. The router network contains $26,464$ nodes with $z=3.26$ and the airport network has $1,573$ nodes with $z=30.39$.
The theoretical predictions and numerical results of the size of the giant NExP component are shown in Fig. 4. We confirm that our theory generates a reliable prediction in the size of the giant NExP component for both the Internet [Fig. 4(a)] and the airport network [Fig. 4(b)]. We also measure the size of the giant connected component for random site percolation (denoted in short as RSP) for comparison callaway ; cohen2000 . We found that $\Phi$ for NExP is higher, compared to the size of giant component by standard node removal scenarios. It concludes that real-world networked systems can be more robust than expected when considering the recovery of exclaves. It is noteworthy that the shape of $\Phi$-curve for the real-world networks is different from those of Fig. 3: It does not display the two-transition character clearly. This is due mainly to the fact that both the real-world networks are “scale-free,” having broad distribution of degrees, driving $q_{c}$ towards zero and $q^{*}$ towards unity. Finite-size effect can also be attributed for the smearing of the transitions in real-world networks. Even so, our theory works remarkably well on these real-world networks.
Figure 4: The size $\Phi$ of the giant NExP component of two real-world networks: (a) the Internet’s router network of autonomous system level and (b) the U.S. airport network. Monte-Carlo simulation results (symbols) averaged over $10^{4}$ independent runs and analytic curves (lines) are shown together.
@@ 5 Summary and discussion
In this study, we have focused on the no-exclaves percolation (NExP) process on random networks and formulated a generating function approach to derive analytical solutions for this model. We identify the percolation threshold and complete transition point for the emergence of the giant NExP component and disappearance of the giant unoccupied component, respectively. In addition, we apply the theory to random regular graphs, Erdős-Rényi graphs, and real-world networks, and confirm the great agreement between our theoretical predictions and Monte Carlo simulations.
From theoretical perspective, our solution offers theoretical understanding for the multiple phase transition structure of NExP on random networks. It shows how the two-threshold behavior associated with the emergence of the giant NExP component and the extinction of the giant unoccupied component emerge as the occupation probability $q$ increases in network NExP. It also suggests that the single discontinuous percolation in 2D NExP nexp may be resolved into two transitions converging at the same point.
From the perspective of network robustness, our study implies that the recovery processes according to the no-exclaves rule can effectively mitigate the impact of random failures of nodes. When we remove nodes with failure probability $1-q$ as random failure problems error_attack ; cohen2000 the giant “occupied” component disappears at $q^{*}$. However, by introducing the no-exclaves rule, we can maintain the giant NExP component at the point $q_{c}$, which is in general less than $q^{*}$. In addition, we also found that the random failure above the complete percolation point $q^{*}$ does not make any effect in the no-exclaves percolation. In this sense, our study suggests an effective way to improve the network robustness with the recovery of failed exclaves.

### v2

@@ 3.4 The average NExP size and percolation threshold
We derive the average size $\chi$ of small NExP components, excluding the giant NExP component. When the giant NExP component does not exist, a node reached by following a randomly chosen link must belong to either a small component with probability $u$ or a giant unoccupied component with probability $1-w$, resulting in the condition $11-w+u=$ 1. It simplifies to the condition $u=w$. With this condition, the average small NExP component size can be expressed as
@@
$=F_{0}(1)+[qG_{0}^{\prime}(1)+(1-q)G_{0}^{\prime}(u)]F_{1}^{\prime}(1),$ (12)
where denotes the derivative with respect to its argument $x$. From Eq. (7), we obtain
$F_{1}^{\prime}(1)=\frac{u}{1-qG_{1}^{\prime}(1)-(1-q)G_{1}^{\prime}(u)}.$ (13)
@@
$\chi=F_{0}(1)+\frac{u[qG_{0}^{\prime}(1)+(1-q)G_{0}^{\prime}(u)]}{[1-qG_{1}^{\prime}(1)-(1-q)G_{1}^{\prime}(u)]}.$ (14)
@@
$q_{c}G_{1}^{\prime}(1)+(1-q_{c})G_{1}^{\prime}(u)=1.$ (15)
The percolation threshold $q_{c}$ of the giant NExP component is in general not amenable to compact expression in terms of $z$ or other moments of $P(k)$ because $w$ and $u$ also contain $q$, yet can be computed numerically. Beyond this threshold $q_{c}$, the size of the giant NExP component continues to grow as more occupied nodes are added.
The NExP threshold $q_{c}$ is smaller than the threshold of the random site percolation $q_{c}^{RSP}$, which is given by $q_{c}^{RSP}G_{1}^{\prime}(1)=1$. The reduction of the threshold is caused by the no-exclaves rule for recovery, which leads to the merging of small and separated occupied components. This contribution is reflected in the second term of Eq. (15). In the region $q_{c}<q<q_{c}^{RSP}$, multiple separated occupied components merge into a giant NExP component during the recovery of exclaves; An example of the merging is depicted in Fig. 1(c).
In NExP, there is another transition where a giant unoccupied component disappears as $q$ increases. In a scenario where a giant unoccupied component vanishes, it becomes possible that the giant NExP component would span the entire nodes within the giant connected component of the underlying undisturbed ($q=1$) network. We refer the point of this transition as the “complete” percolation point, denoted as $q^{*}$. The complete percolation point $q^{*}$ corresponds to the location in which the size $\Phi$ of the giant NExP component becomes unity if all nodes of the underlying network belong to a single connected component. The location of $q^{*}$ can be identified by the divergence of the average size $\langle s\rangle$ of unoccupied components. The condition for $q^{*}$ is given from Eq. (3) as
@@
which is directly related to the well-known result for the percolation on random networks [11, 12, 35].
@@
The transitions can be interpreted by the stability analysis of the solutions of Eqs. (5) and (9). The complete percolation point $q^{*}$ corresponds to the point where the trivial fixed point of $w$ in Eq. (5) becomes unstable. The trivial solution of Eq. (5) is $w=1$ since $G_{1}(1)$ is normalized to be $1$. The solution $w=1$ corresponds to no giant unoccupied component considering the meaning of $w$. It means that the giant unoccupied component disappears at $q^{*}$. On the other hand, the NExP threshold $q_{c}$ can be obtained by the stability analysis of Eq. (9). The trivial solution of Eq. (9) is $u=w$ because Eq. (9) reduces into Eq. (5) when $u=w$, or equivalently $11-w+u=$ 1. The solution $u=w$ represents no giant NExP component and it becomes unstable at $q_{c}$. Therefore another solution becomes stable, leading to the emergence of the giant NExP component.
@@
These two transition points $q_{c}$ and $q^{*}$ do not in general coincide on random networks. The percolation transition $q_{c}$ is generally smaller than the complete point $q^{*}$. Therefore, there can be three phases in NExP, depending on the existence and extinction of the giant NExP component and the giant unoccupied component: We have phases in which i) the giant unoccupied component exists but giant NExP component does not when $q<q_{c}$; ii) both the giant NExP component and the giant unoccupied component coexist when $q_{c}<q<q^{*}$; iii) the giant NExP component exists but giant unoccupied component does not when $q>q^{*}$. The multiple percolation transitions in network NExP are in stark contrast to the 2D NExP in which the two occur at the same point [26]. Our solutions suggest that the discontinuous percolation transition into complete percolation displayed by 2D NExP can be resolved into two transitions combined into one.
3.5 Critical exponents of NExP
We examine the critical exponents of NExP in the vicinity of the NExP threshold $q_{c}$ by using generating function method [36]. We first determine the order parameter critical exponent $\beta$, indicating that $\Phi\sim(q-q_{c})^{\beta}$. The singular behavior of $\Phi$ stems from the self-consistency equations, Eqs. (5) and (9). We examine Eq. (5) by rewriting with $w=w_{c}+\epsilon_{w}$ and $q=q_{c}+\delta$. Then, Eq. (5) becomes $w_{c}+\epsilon_{w}=(q_{c}+\delta)+(1-q_{c}-\delta)G_{1}(w_{c}+\epsilon_{w})$. Keeping only the leading order of $\epsilon_{w}$ for a small $\epsilon_{w}$, Eq. (5) implies
$\epsilon_{w}$ $\sim\frac{1-G_{1}(w_{c})}{q_{c}G_{1}^{\prime}(1)}\delta.$ (18)
Similarly, we expand Eq. (9) for $q=q_{c}+\delta$, $w=w_{c}+\epsilon_{w}$. and $u=u_{c}-\epsilon_{u}$ and obtain the relation
$\epsilon_{u}\sim\frac{2[G_{1}^{\prime}(1)-G_{1}^{\prime}(u_{c})]}{q_{c}G_{1}^{\prime\prime}(1)+(1-q_{c})G_{1}^{\prime\prime}(u_{c})}\delta.$ (19)
Finally, the size of the giant NExP component grows near the NExP threshold as:
$\Phi\sim c_{w}\epsilon_{w}+c_{u}\epsilon_{u}\sim(q-q_{c})^{\beta},$ (20)
leading to $\beta=1$ which agrees with the mean-field percolation result.
By employing the ansatz $Q(\phi)\sim\phi^{1-\tau}$ at $q=q_{c}$, we can derive the exponent $\tau$. When we expand Eq. (7) at $q=q_{c}$ for $x=1-\epsilon$ and $F_{1}(x)=u_{c}-\Delta$, the lowest order yields
$\epsilon\sim\frac{q_{c}G_{1}^{\prime\prime}(1)+(1-q_{c})G_{1}^{\prime\prime}(u_{c})}{2u_{c}}\Delta^{2}.$ (21)
At the NExP threshold $q_{c}$, the average size of $\phi$ diverges, leading to the fact $2<\tau\leq 3$. Then the generating function $F_{1}(x)$ should satisfy $u_{c}-F_{1}(1-\epsilon)\sim\epsilon^{\tau-2}$. Given that $\Delta\sim\epsilon^{1/2}$, we arrives at $\tau=5/2$, which corresponds to the same value of the mean-field percolation class.
Other critical exponents are determined by scaling relations [1, 36], and thus we suggest that the NExP belongs to the mean-field universality class, based on the fact that $\beta=1$ and $\tau=5/2$. In the case of scale-free networks, the critical exponents would differ from the mean-field values due to the dependency on degree exponents. However, we have not considered the effects of degree heterogeneity of scale-free networks in this study.
4 Specific Results
@@ 4.1 On Random Regular Networks
When $w=1$ for $q>q^{*}$, $u$ becomes 00. If $q<q^{*}$ indicating that $w=q/(1-q)$, the possible solutions of $u$ are either $u=q/(1-q)$ and $(1-2q)^{2}/(1-q)$. The solution $u=w=q/(1-q)$ corresponds to the case when there is no giant NExP component. If $u=(1-2q)^{2}/(1-q)$, $1-w+u$ becomes less than unity meaning that there is a giant NExP component. The NExP threshold between two phases is given by
$w(q_{c})$ $=\frac{1}{z}=\frac{1}{3},$ (28)
which leads to
@@
Therefore we have three regimes of no-exclaves percolation in RR networks: i) giant NExP component does not exist but the giant unoccupied component exists ($q<q_{c}$); ii) both the giant NExP component and the giant unoccupied component coexist ($q_{c}<q<q^{*}$); and iii) only the giant NExP component whose size is unity exists ($q>q^{*}$). Since we can express $u$ and $w$ in closed forms for RR networks with $z=3$, we can explicitly obtain the expression for the giant NExP component size $\Phi$ as:
@@
As shown in Fig. 2, the probabilities $w$ and $u$, and the giant NExP component size $\Phi$ as a function of the occupation probability $q$ show perfect agreements with the Monte Carlo simulation results.
Figure 2: Analytic predictions of $\Phi$, $w$, and $1-w+u$ with respect to $q$ on RR networks with $z=3$ and $N=10^{6}$ are shown. Theoretical results (lines) and simulation results (points) are shown together. Error-bars denote the standard deviation from $10^{6}$ independent runs.
@@ 4.2 On Erdős-Rényi Networks
We consider the NExP model on Erdős-Rényi (ER) networks, which have a Poisson degree distribution $P(k)=e^{-z}z^{k}/k!$ in the large $N$ limit, where $z$ is the average degree. The generating functions for ER networks are given by $G_{0}(x)=e^{z(x-1)}$ and $G_{1}(x)=e^{z(x-1)}$. This leads us to the coupled equations for the variables $w$ and $u$, given by
@@
The complete percolation point and the percolation threshold for NExP can be determined by Eqs. (17) and (15), respectively. These conditions show the instability of the trivial solutions, $w=1$ for the complete percolation point and $u=w$ for the percolation threshold, respectively. For ER networks, the specific values for these thresholds are given by
$q^{*}=1-\frac{1}{z}\quad\textrm{and}\quad w(q_{c})=\frac{1}{z}.$ (40)
We have three regimes in the NExP model, similar to RR networks. i) When $q<q_{c}$, there is no giant NExP component but the giant unoccupied component exists. ii) When $q_{c}<q<q^{*}$, both the giant NExP component and the giant unoccupied component coexist. iii) When $q>q^{*}$, only the giant NExP component exists. Furthermore, when $q>q_{c}$, the giant NExP component size $\Phi$ can be expressed as
$\Phi$ $=1-qe^{z(u-w)}-(1-q)[1-e^{z(w-1)}+e^{z(u-1)}],$ (41)
where $w$ and $u$ can be obtained by Eqs. (38, 39).
We compare the theoretical results derived using the generating function method with the Monte Carlo simulations on ER networks of average degree $z=3/2,2,3$ and network size $N=10^{6}$. Figure 3 shows the giant NExP component size $\Phi$ obtained from our theory (lines) and from the numerical simulations (points) as a function of $q$. We observe excellent agreements between our theoretical predictions and the simulation results. Furthermore, we validate the locations of the percolation transition, $q_{c}$, and the complete transition point, $q^{*}$, which were predicted from Eq. (40).
Figure 3: The giant NExP component size $\Phi$ as a function of the occupation probability $q$ on ER networks with mean degree $z=3/2,2$, and $3$ and $N=10^{6}$. Numerical results (points) averaged over $10^{6}$ independent runs and analytic curves (lines) are shown together.
@@ 4.3 On real-world networks
We test our theory for NExP on top of real-world networks built from empirical data. We use the two datasets: the Internet of an autonomous system level compiled by CAIDA [37] and the United States airport network obtained from OpenFlights [38]. The former represents the router network of Internet reconstructed from the packet flows between neighboring peers at the autonomous system level, collected on November 5, 2007. The U.S. airport network represents the connections from the origin to destination airports in the year 2010. For the sake of simplicity, we ignore directionality and weights of the links from the data when we construct the networks. The router network contains $N=26,464$ nodes with the mean degree $z=3.26$ and the airport network has $1,573$ nodes with $z=30.39$.
The theoretical predictions and numerical simulation results for the giant NExP component size $\Phi$ are shown in Fig. 4. We confirm that our theory generates a reliable prediction in the size of the giant NExP component for both the Internet [Fig. 4(a)] and the airport network [Fig. 4(b)]. We also measure the size of the giant connected component for random site percolation (denoted in short as RSP) for comparison [11, 12]. We found that $\Phi$ for NExP is higher, compared to that for the standard node removal scenarios. It concludes that real-world networked systems can be more robust than expected when considering the recovery of exclaves. It is noteworthy that the shape of $\Phi$-curve for the real-world networks is different from those of Fig. 3: It does not display the two-transition character clearly. This is due mainly to the fact that both the real-world networks are “scale-free,” having broad distribution of degrees, driving $q_{c}$ towards zero and $q^{*}$ towards unity. Finite-size effects can also be attributed for the smearing of the transitions in real-world networks. Even so, our theory works remarkably well on these real-world networks.
Figure 4: The size $\Phi$ of the giant NExP component of two real-world networks: (a) the Internet’s router network at the autonomous system level and (b) the U.S. airport network. Monte-Carlo simulation results (symbols) averaged over $10^{4}$ independent runs and analytic curves (lines) are shown together.
@@ 5 Summary and discussion
In this study, we have focused on the no-exclaves percolation (NExP) process on random networks and formulated a generating function approach to derive analytical solutions for this model. We identify the percolation threshold and the complete transition point for the emergence of the giant NExP component and the disappearance of the giant unoccupied component, respectively. In addition, we apply the theory to random regular graphs, Erdős-Rényi graphs, and real-world networks, and confirm great agreements between our theoretical predictions and Monte Carlo simulations.
From theoretical perspective, our solution offers theoretical understanding for the multiple phase transition structure of NExP on random networks. It shows how the two-threshold behavior associated with the emergence of the giant NExP component and the extinction of the giant unoccupied component emerge as the occupation probability $q$ increases in network NExP. It also suggests that the single discontinuous percolation in 2D NExP [26] may be resolved into two transitions converging at the same point.
From the perspective of network robustness, our study implies that the recovery processes according to the no-exclaves rule can effectively mitigate the impact of random failures of nodes. When we remove nodes with failure probability $1-q$ as random failure problems [2, 12], the giant “occupied” component disappears at $q^{*}$. However, by introducing the no-exclaves rule, we can maintain the giant NExP component at the point $q_{c}$, which is in general less than $q^{*}$. In addition, we also found that the random failure above the complete percolation point $q^{*}$ does not make any effect in the no-exclaves percolation. In this sense, our study suggests an effective way to improve the network robustness with the recovery of failed exclaves.